OpenStudy (anonymous):

Find b and c so that (5,b,c) is orthogonal to both (1,2,3) and (1,-2,1)

OpenStudy (helder_edwin):

let \(v=(5,b,c)\) and \(w_1=(1,2,3)\) and \(w_2=(1,-2,1)\) solve the homogeneus linear system \[ \large \begin{cases} v\cdot w_1=0\\ v\cdot w_2=0 \end{cases} \]

OpenStudy (anonymous):

I did that and I didn't get the right answer.

OpenStudy (helder_edwin):

your system would be \[ \large \begin{cases} 5+2b+3c=0\\ 5-2b+c=0 \end{cases} \]

OpenStudy (helder_edwin):

adding both equations u get \[ \large 10+4c=0\Rightarrow c=-10/4=-5/2 \]

OpenStudy (helder_edwin):

then \[ \large 5-2b+(-5/2)=0\Rightarrow b=(5/2-5)/(-2)=5/2-5/4=5/4 \]

OpenStudy (anonymous):

Thank you!

OpenStudy (helder_edwin):

u r welcome